OpenStudy (anonymous):

Find the polar equation that has the same graph as the equation in x and y: x = -3

OpenStudy (anonymous):

\[x=r \cos \theta=-3=3*-1,r=3,\cos \theta=-1=\cos \pi,\theta=\pi \] polar co-ordinates are (3,pi )

OpenStudy (jdoe0001):

x = -3 => x+3 = 0 which will give the ordered pair of (-3, 0) whilst x = -3, y = 0 \(\bf r = \sqrt{(-3)^2+0^2} \implies r = 3\\ tan\left(\frac{0}{-3}\right) = 0 \implies tan^{-1}(0) = \pi\)

OpenStudy (anonymous):

According to the back of my book (because it is an odd problem), the answer is \[r = -3 \sec \theta\]

OpenStudy (anonymous):

I just don't know how to get that..

OpenStudy (jdoe0001):

in all honesty, I can see r = -3 and +3 while the angle being 0 or pi and both values will yield a -3

OpenStudy (anonymous):

\[r=3 \cos \pi \]

OpenStudy (anonymous):

r is always positive

OpenStudy (anonymous):

@jdoe0001 okay, so what about the sec (theta)

OpenStudy (jdoe0001):

\(\bf x = rcos(\theta) \textit{ thus then}\\ x = -3 \implies rcos(\theta) = -3 \implies r = \cfrac{-3}{cos(\theta)}\\ r = \cfrac{-3}{1} \cdot \cfrac{1}{cos(\theta)}\)

OpenStudy (anonymous):

Great!! Thank you!! @jdoe0001